A skier is pulled by a tow rope up a frictionless ski slope that makes an angle of 8.9° with the horizontal. The rope moves parallel to the slope with a constant speed of 1.1 m/s. The force of the rope does 640 J of work on the skier as the skier moves a distance of 5.9 m up the incline. (a) If the rope moved with a constant speed of 2.4 m/s, how much work would the force of the rope do on the skier as the skier moved a distance of 5.9 m up the incline? At what rate is the force of the rope doing work on the skier when the rope moves with a speed of (b) 1.1 m/s and (c) 2.4 m/s?
Answer #1

a)

force applied by rope

T = W/d = 640 / 5.9 = 108.475 N

=====

b)

P = T * v = 108.475* 1.1 = 119.322 W

c)

P = 108.475* 2.4 = 260.34 W

=====